Linkin Corporation is considering purchasing a new delivery truck. The truck has many advantages over the company’s current truck (not the least of which is that it runs). The new truck would cost $57,500. Because of the increased capacity, reduced maintenance costs, and increased fuel economy, the new truck is expected to generate cost savings of $8,300. At the end of 8 years the company will sell the truck for an estimated $28,800. Traditionally the company has used a rule of thumb that a proposal should not be accepted unless it has a payback period that is less than 50% of the asset’s estimated useful life. Larry Newton, a new manager, has suggested that the company should not rely solely on the payback approach, but should also employ the net present value method when evaluating new projects. The company’s cost of capital is 8%. Click here to view PV table. (a) Compute the cash payback period and net present value of the proposed investment. (If the net present value is negative, use either a negative sign preceding the number eg -45 or parentheses eg (45). Round answer for present value to 0 decimal places, e.g. 125. Round answer for Payback period to 1 decimal place, e.g. 10.5. For calculation purposes, use 5 decimal places as displayed in the factor table provided.)
Answer #1
Investment cost 57500
Divide by Annual cash flows 8300
Cash payback period 6.9 years
Present value of annual cash flows 47697 =8300*5.74664
Present value of salvage value 15560 =28800*0.54027
Total Present value 63257
Less: Investment cost 57500
Net present value 5757
PV factor computation:
(1-(1.08)^-8)/0.08 = 5.74664
1/1.08^8 = 0.54027